OpenStudy (anonymous):

what are the real or imaginary solutions of the polynomial equation x^3-216=0

OpenStudy (ranga):

Write 216 as 6^3 and use the following identity to factor: \[\Large a ^{3} - b ^{3} = (a - b)(a ^{2} + ab + b ^{2})\]

Directrix (directrix):

x^3-216=0 (x - 6) (x² + 6x + 36) = 0 Apply Zero Product Property x - 6 = 0 or (x² + 6x + 36) = 0 Hint: Use the Quadratic Formula to solve for x in (x² + 6x + 36) = 0 @steph19745